# Imitation Learning for Combinatorial Optimisation under Uncertainty

### v1

@@ 3.3.2 End-to-End Learning
End-to-end learning refers to a modelling paradigm that tightly integrates machine learning and optimisation within a single framework. In this setting, the learning model either directly predicts solutions to optimisation problems, corresponding to the AIT, or produces partially specified optimisation models whose parameters are inferred from raw or structured input data, corresponding to the PPT.
@@
Action Imitation
An early attempt to directly predict solutions to difficult combinatorial problems, such as the Travelling Salesman Problem (TSP) is due to [vinyals2015pointer], who propose pointer networks trained in a supervised manner on solved instances. Their approach demonstrates limited generalisation to instances of varying sizes. Building on this idea, [li2018combinatorial] and [joshi2019efficient] develop supervised models for static combinatorial optimisation problems.
@@
ML-augmented optimisation and end-to-end learning thus represent two dominant paradigms for combining machine learning with optimisation. While ML-augmented approaches embed learning within specific algorithmic components such as branching, cutting, or search space reduction, end-to-end methods aim to predict decisions or optimisation parameters directly through unified learning pipelines. Across both paradigms, supervision plays a central role, yet the nature of the expert varies substantially, ranging from full-information solutions and heuristics to intermediate solver states. These choices have significant implications for learning performance, generalisability, and computational cost. Understanding how experts are defined and utilised is therefore crucial for algorithm design. We next introduce a taxonomy of expert types to clarify these distinctions and to guide the development of imitation learning methods for sequential decision problems.
@@ 4 Expert Taxonomy
An expert in imitation learning is a policy, or more generally a decision-making function, that specifies the desired or target behaviour in a sequential decision problem. The expert guides the learning agent by demonstrating actions in given states, typically under the assumption of access to an optimal or near-optimal strategy with respect to a known or implicit objective. Alternative terms used in the literature include oracle, demonstrator, optimal policy, ground truth, reference policy, and target. Broadly, we distinguish between two categories of experts.
  • Human Expert. Human experts rely on practical experience and contextual understanding to make decisions under uncertainty. In dynamic environments such as vehicle dispatching or healthcare scheduling, they implicitly account for real-time information, operational constraints, and institutional knowledge. For example, [ozarik2024machine] study last-mile delivery and imitate driver decisions that capture practical insights not explicitly modelled in traditional optimisation frameworks. While such decisions often reflect nuanced judgement, they may also be inconsistent or influenced by cognitive biases. Moreover, collecting sufficiently large and high-quality datasets from human experts can be challenging, particularly in domains where expert time is scarce or costly.
  • Algorithmic Expert. Algorithmic experts generate decisions using mathematical models, optimisation solvers, or rule-based and heuristic procedures. These experts can typically produce large volumes of consistent and high-quality demonstrations. Their structure depends on how the decision objective is defined and how uncertainty, denoted by $\boldsymbol{\xi}$, is treated. The vector $\boldsymbol{\xi}$ represents exogenous factors such as demand, travel times, or arrival processes, which are not fully known at the time decisions are made.
@@
Ideally, for a sequential decision problem, the optimal action at decision epoch $k$ is obtained by solving
$a_{k}^{*}=\operatorname*{arg\,min}_{a_{k}}\mathbb{E}\Big\{C_{k}(x_{k},a_{k})+V_{k+1}(x_{k+1})\,\big|\,x_{k}\Big\},$ (11)
where the expectation is taken with respect to future uncertainty. For many real-world business applications, computing such optimal actions, even offline, is often impractical due to problem complexity. Consequently, it is common in operations research to design algorithms that produce near-optimal solutions—decisions that are not globally optimal but are sufficiently close to be useful in practice. When such solutions can be computed within reasonable time limits, they can serve as expert demonstrations for IL.
Despite the widespread use of approximate or surrogate experts, the literature lacks a clear and systematic taxonomy for defining experts and their roles within IL. To address this gap, we propose a structured taxonomy of expert types to guide the design and application of IL methods in sequential decision problems.
@@ 4.1 Expert Decision Models under Uncertainty
We further classify algorithmic experts according to how they handle uncertainty $\boldsymbol{\xi}$, which captures factors such as future demand, travel times, or patient arrivals that are not fully observed at the time of decision making. The treatment of $\boldsymbol{\xi}$ determines whether an expert behaves myopically or anticipates future outcomes, and whether it is relatively simple or highly adaptive. We outline five classes of algorithmic experts, distinguished by how uncertainty is incorporated into their decision models.
@@
    Two-Stage Stochastic Expert - In the two-stage stochastic framework, uncertainty $\boldsymbol{\xi}$ is not observed at the initial decision stage. The expert first selects a here-and-now action and subsequently adapts through recourse decisions once uncertainty is revealed. This structure captures limited anticipation while retaining tractability and is widely used in stochastic programming.
@@ 4.2 Expert Optimality
Even when employing the expert types defined above, the underlying optimisation problem may remain difficult to solve. Many deterministic optimisation problems are $\mathcal{NP}$-hard, rendering exact solution approaches computationally impractical in many real-world settings. In such cases, practitioners rely on heuristics or impose stopping criteria when using optimisation solvers. We therefore classify experts according to the optimality guarantees of the models they solve.
  • Task-Optimal Expert. A task-optimal expert solves its underlying optimisation model to optimality, assuming the problem is tractable and can be solved exactly within reasonable computational limits. Such experts are most applicable when the problem structure permits efficient solution methods. For example, a strong branching expert evaluates linear programming relaxations corresponding to all fractional variables and selects the branching variable that yields the greatest improvement in the objective [Nair2020SolvingMI, alvarez2016online]. Optimal solutions may be obtained using a variety of specialised solvers, including mixed-integer programming solvers [sun2022learning], SAT solvers [sun2024enhancing], and dedicated solvers such as Concorde for the travelling salesman problem [sun2021generalization]. We denote the policy of a task-optimal expert by $\pi^{*\cdot}$ and the corresponding optimal action by $a^{*\cdot}$ where $\cdot$ would denote the superscript of the respective experts from Table˜1.
  • Approximate Expert. An approximate expert produces actions that aim to closely approximate optimal or near-optimal decisions while trading off solution quality for computational efficiency. Several commonly used approximation strategies include:
    • Time limits - The solver is terminated after a fixed time budget, returning the best solution found so far [li2025learning, baty2024combinatorial, pham2023prediction, sun2021generalization].
    • Optimality gaps -The solver terminates once a predefined optimality gap is achieved, ensuring a bounded deviation from optimality [larsen2018predicting].
    • Objective-based refinement - Alternative formulations or local improvements are used to incrementally improve a solution relative to the objective function [ding2020accelerating].
    • Heuristics - High-quality heuristic methods may be used directly to generate solutions. For example, [baty2024combinatorial] employ hybrid genetic search for the prize-collecting TSP. Alternatively, a finite set of candidate actions may be evaluated, and the best-performing action selected [song2020general].
@@ 4.4 Labelling strategy
      Single call - A single interaction with the expert yields multiple candidate actions for a given state. This occurs, for example, when solving an integer programme using a mathematical programming solver that returns multiple feasible solutions. In [khalil2022mip], CPLEX is allowed to run for 60 minutes to generate a solution pool for each instance, terminating after identifying 1,000 solutions with objective values within 10% of the best solution found.
@@ 4.5 Number of Experts Used
  • Single Expert- The conventional setting in imitation learning, where a single, typically optimal expert provides guidance. The learner assumes the expert is both reliable and inexpensive to query, consistently supplying accurate actions for all states. This simplifies learning but may limit adaptability in complex or diverse environments. Most literature uses a single expert to build the dataset.
  • Multi-Expert- In some cases, multiple experts with complementary specialisations are available. Their actions can be aggregated into a composite "meta-expert", combining diverse expertise to enhance robustness and flexibility across different aspects of the task. [song2020general] uses two different experts to get demonstrations. First, they use an expensive expert to gather initial expert demonstrations, which are used to train initial imitation learning policies. Next, a simpler retrospective expert is used to augment the dataset. [rautenstrauss2025optimization] uses a set of non-optimal policies to label states for which the true hindsight expert is expensive. These states generally represent initial decision epochs in the horizon. [sun2023mega] introduces MEGA-DAgger, a variant of DAgger designed for interactive learning with multiple imperfect experts.
@@ 4.6 Trajectory Control
Moreover, to reduce reliance on costly expert calls, the expert is queried only for critical cases, such as novel or poorly handled states. This selective querying improves sample efficiency and focuses learning on high-impact scenarios. [menda2019ensembledagger], [hoque2021lazydagger] describe the Ensemble DAgger and Lazy DAgger algorithms, where the expert is selectively queried based on the discrepancy measure between the expert and learned policy actions. However, such variants of the DAgger algorithms have not been experimented with for SDP in operations research.
@@
Using the expert taxonomy, we now propose and test a general DAgger-based algorithm that links each dimension of the taxonomy to a corresponding design choice in the learning procedure, enabling a structured comparison of expert configurations.
5 Problem Description
The classical Generalised Assignment Problem (GAP) has been extensively studied since the seminal work of [balachandran1976integer]. The problem consists of assigning a set of items (or jobs) to a set of agents (or machines) such that each job is assigned to exactly one agent, agent capacity constraints are respected, and the total assignment cost or profit is optimised. GAP has found applications in a wide range of domains, including job scheduling [balachandran1976integer], routing [fischetti2019machine], and patient admission scheduling [ceschia2011local]. A standard integer programming formulation of the GAP is given by
$\max\quad$ $\sum_{i=1}^{m}\sum_{j=1}^{n}c_{ij}x_{ij}$ (12)
s.t. $\sum_{i=1}^{m}x_{ij}=1,\quad$ $\forall j=1,\dots,n,$ (13)
$\sum_{j=1}^{n}d_{ij}x_{ij}\leq L_{i},\quad$ $\forall i=1,\dots,m,$ (14)
$x_{ij}\in\{0,1\},\quad$ $\forall i=1,\dots,m,\ \forall j=1,\dots,n.$ (15)
Here, $c_{ij}\in\mathbb{R}_{+}$ denotes the profit obtained by assigning job $j$ to agent $i$, $d_{ij}\in\mathbb{R}_{+}$ is the resource consumption of agent $i$ when processing job $j$, $L_{i}\in\mathbb{R}_{+}$ is the capacity of agent $i$, and $x_{ij}$ is a binary decision variable equal to $1$ if job $j$ is assigned to agent $i$ and 0 otherwise.
5.1 Physician-to-Patient Assignment
Appointment requests arrive independently over time and are handled individually by a receptionist. At each arrival, the receptionist must decide whether to accept or reject the request based on the current system state. If the request is accepted, the receptionist must assign the patient to a physician. Each patient $k$ has a preferred physician $p_{k}$ among a set of eligible physicians $\mathcal{P}_{k}\subseteq\mathcal{P}$. Each patient is characterised by a service duration $d_{k}\in\mathbb{R}_{+}$, which is estimated at the time of the call, and a priority class $r_{k}\in\{1,2\}$. The arrival process is such that call from priority 1 patients arrive near the end of the horizon.
Rejecting a patient of priority $r_{k}$ incurs a penalty $c^{\text{rej}}_{r_{k}}$, with $c^{\text{rej}}_{1}>c^{\text{rej}}_{2}$. In addition, assigning a patient to an undesired but eligible physician incurs a preference penalty $c^{\text{pref}}_{r_{k}}$. The primary source of uncertainty in the system arises from the random number and sequence of patient arrivals within a session. This creates a trade-off between accepting lower-priority patients early and preserving capacity for potential high-priority patients who may arrive later. Furthermore, assigning patients to undesired physicians may reduce future capacity, potentially forcing rejections later in the booking horizon.
The resulting problem is a sequential decision problem in which each decision epoch corresponds to a patient arrival, and the action is to assign the patient to an eligible physician or to reject the request. The offline mixed-integer programming formulation, denoted by $(\textbf{{MIP}}_{\textbf{PPA}})$, is provided in the appendix.
5.2 MDP Formulation
We now formulate the PPA problem as a Markov decision process (MDP).
State Variable – At decision epoch $k$, the system state is given by
$x_{k}=\big(x^{k}_{\text{patient}},x^{k}_{\text{physician}},x^{k}_{\text{eligibility}}\big).$
The component $x^{k}_{\text{patient}}=[d_{k},r_{k}]$ represents the characteristics of the current patient. The component
$x^{k}_{\text{physician}}=[x^{k}_{p}]_{p\in\mathcal{P}}$
captures the status of all physicians, where
$x^{k}_{p}=\big(\mathbbm{1}_{p=p_{k}},x^{k}_{l},x^{k}_{p1},x^{k}_{pw}\big).$
Here, $\mathbbm{1}_{p=p_{k}}$ indicates whether physician $p$ is the patient preferred physician, $x^{k}_{l}$ denotes the remaining assignment capacity, $x^{k}_{p1}$ is the number of priority-1 patients already assigned, and $x^{k}_{pw}$ is the current workload. The eligibility vector is given by a list of indicator functions
$x^{k}_{\text{eligibility}}=[\mathbbm{1}_{p\in\mathcal{P}_{k}}]_{p\in\mathcal{P}}.$
Decision Variable – At each decision epoch $k$, the receptionist selects an action
$a_{k}\in\mathcal{A}_{k}=\{0\}\cup\mathcal{P}_{k},$
where $a_{k}=0$ corresponds to rejecting the request and $a_{k}=p$ corresponds to assigning the patient to physician $p$. A physician $p$ is eligible only if $d_{k}+x^{k}_{pw}\leq L_{p}$, ensuring feasibility of all actions in $\mathcal{A}_{k}$.
Exogenous Information – The exogenous information at epoch $k$ corresponds to the characteristics of the next arriving patient and is denoted by
$\xi_{k}=[d_{k+1},r_{k+1},p_{k+1},\mathcal{P}_{k+1}].$
Transition Function – The transition function $X^{M}(\cdot)$ maps the system from state $x_{k}$ to $x_{k+1}$ according to
$x_{k+1}=X^{M}(x_{k},a_{k},\xi_{k}).$
Specifically, for the selected physician $p$,
  • $x^{k+1}_{l}=x^{k}_{l}-\mathbbm{1}_{a_{k}=p}$,
  • $x^{k+1}_{p1}=x^{k}_{p1}+\mathbbm{1}_{a_{k}=p}\mathbbm{1}_{r_{k}=1}$,
  • $x^{k+1}_{pw}=x^{k}_{pw}+\mathbbm{1}_{a_{k}=p}d_{k}$.
$C_{k}(x_{k},a_{k})=\mathbbm{1}_{a_{k}=0}\,c^{\text{rej}}_{r_{k}}+\mathbbm{1}_{a_{k}\neq 0}\,\mathbbm{1}_{a_{k}\neq p_{k}}\,c^{\text{pref}}_{r_{k}}.$
@@ 6 Algorithm
In this section, we present a general-purpose algorithm for learning from experts in sequential decision problems. The proposed algorithm, shown in Algorithm˜1, explicitly incorporates the expert taxonomy introduced in Section˜4. A summary of the algorithm parameters is provided in Table˜2.
The algorithm is based on the Dataset Aggregation (DAgger) framework, which follows an iterative policy learning paradigm inspired by online learning and assumes access to one or more experts. Let $\mathcal{E}$ denote the set of available experts. At each learning iteration, indexed by $i$, the learner retrains its policy using all states encountered so far, together with the corresponding expert-labelled actions. This iterative process allows the learner to correct earlier mistakes and progressively improve their performance.
The procedure starts from an initial policy $\pi_{0}$, which may be obtained from expert demonstrations or derived from a heuristic. This policy is then executed for $H$ simulated episodes using the decision rule, each consisting of $K$ decision epochs. At each epoch, the learner observes the current state and may query an expert up to $J$ times, as determined by the function $F_{j}$. Each query returns an expert action, potentially from different experts in $\mathcal{E}$, depending on the expert-selection function $F_{e}$.
We denote an expert action by $a^{\cdot}_{ijk}$, where $i$ indexes the learning iteration, $j$ the expert call within the epoch, and $k$ the decision epoch. For clarity, we omit the episode index in this notation. This may be task-optimal or approximate depending on the complexity of the underlying optimisation problem. The parameter $\kappa$ specifies the list of expert stopping criteria, thereby controlling the quality of the returned action. A decision rule (DR) then determines which action $a_{k}$ is to be executed in the environment, influencing the trajectory of the episode. The resulting state–action pairs are stored in the dataset before the transition to the next state occurs.
After completing $H$ episodes, the dataset is augmented with the newly collected expert-labelled samples, and a new policy $\pi_{i+1}$ is trained. This process is repeated until a stopping condition is satisfied, yielding progressively refined policies. The following paragraphs relate the components of the expert taxonomy to Algorithm˜1.
@@
Table 2: List of parameters for Algorithm˜1
@@
4   Initialize dataset $\mathcal{D}\leftarrow\emptyset$, initial policy $\pi_{0}$, and iteration counter $i\leftarrow 0$;
5

### v2

@@ 3.3.2 End-to-End Learning
- End-to-end learning refers to a modelling paradigm that tightly integrates ML and optimisation within a single framework. In this setting, the learning model either directly predicts solutions to optimisation problems, corresponding to the AIT, or produces partially specified optimisation models whose parameters are inferred from raw or structured input data, corresponding to the PPT.
@@
Action Imitation Task
- An early attempt to directly predict solutions to difficult combinatorial problems, such as the Travelling Salesman Problem (TSP), is due to [vinyals2015pointer], who propose pointer networks trained in a supervised manner on solved instances. Their approach demonstrates limited generalisation to instances of varying sizes. Building on this idea, [li2018combinatorial] and [joshi2019efficient] develop supervised models for static CO problems.
@@
ML-augmented optimisation and end-to-end learning thus represent two dominant paradigms for combining ML with optimisation. While ML-augmented approaches embed learning within specific algorithmic components such as branching, cutting, or search space reduction, end-to-end methods aim to predict decisions or optimisation parameters directly through unified learning pipelines. Across both paradigms, supervision plays a central role, yet the nature of the expert varies substantially, ranging from full-information solutions and heuristics to intermediate solver states. These choices have significant implications for learning performance, generalisability, and computational cost. Understanding how experts are defined and utilised is, therefore, crucial for algorithm design. We next introduce a taxonomy of expert types to clarify these distinctions and to guide the development of IL methods for SDPs.
@@ 4 Expert Taxonomy
In IL, an expert is a policy, or more generally a decision-making function, that specifies the desired behaviour in a sequential decision problem. The expert guides the learning agent by providing action demonstrations for given states, typically assuming access to an optimal or near-optimal strategy under a known or implicit objective. Related terms used in the literature include oracle, demonstrator, optimal policy, ground truth, reference policy, and target. We broadly distinguish between two categories of experts.
  • Human Expert – Human experts rely on experience and contextual understanding to make decisions under uncertainty. In dynamic settings such as vehicle dispatching or healthcare scheduling, they implicitly incorporate real-time information, operational constraints, and institutional knowledge. For example, [ozarik2024machine] study last-mile delivery by imitating driver decisions that capture practical insights not explicitly represented in optimisation models. While human decisions can be nuanced, they may also be inconsistent or subject to cognitive bias. Moreover, collecting large and reliable datasets from human experts is often difficult due to limited availability and high cost.
  • Algorithmic Expert – Algorithmic experts generate decisions using mathematical models, optimisation solvers, or rule-based and heuristic methods. Such experts can typically produce large volumes of consistent demonstrations. Their design depends on how the decision objective is specified and how uncertainty, denoted by $\boldsymbol{\xi}$, is handled. The vector $\boldsymbol{\xi}$ captures exogenous factors such as demand, travel times, or arrival processes that are not fully known at the time of decision making.
@@
Ideally, for an SDP, the optimal action at decision epoch $k$ is obtained by solving (7) where the expectation is taken with respect to future uncertainty.
$a_{k}^{*}=\operatorname*{arg\,min}_{a_{k}}\mathbb{E}_{\xi}\Big\{C_{k}(x_{k},a_{k})+V_{k+1}(x_{k+1})\,\big|\,x_{k}\Big\},$ (7)
For many real-world business applications, computing such optimal actions, even offline, is often impractical due to problem complexity. Consequently, it is common in operations research to design algorithms that produce near-optimal solutions—decisions that are not globally optimal but are sufficiently close to be useful in practice. When such solutions can be computed within reasonable time limits, they can serve as expert demonstrations for IL.
Despite the widespread use of approximate or surrogate experts, the literature lacks a clear and systematic taxonomy for defining experts and their roles within IL. To address this gap, we propose a structured taxonomy of expert types to guide the design and application of IL methods in SDPs.
@@ 4.1 Expert Decision Models under Uncertainty
We further classify algorithmic experts according to how they handle uncertainty $\boldsymbol{\xi}$ at the time of decision making. The treatment of $\boldsymbol{\xi}$ determines whether an expert behaves myopically or anticipates future outcomes, and whether it is relatively simple or highly adaptive. We outline five classes of algorithmic experts, distinguished by how uncertainty is incorporated into their decision models.
@@
    Two-Stage Stochastic Expert - In the two-stage stochastic framework, uncertainty $\boldsymbol{\xi}$ is not observed at the initial decision stage. The expert first selects a here-and-now (first stage) action and subsequently adapts through recourse decisions (second stage) once uncertainty is revealed. The cost function can be divided into two parts, where the first part $Q_{1}(a)$ represents the first stage cost and $Q_{2}(a,\xi)$ denotes the second stage cost. This structure captures limited anticipation by committing to an initial action before uncertainty is known and then optimally adjusting with recourse actions after the uncertainty is revealed. Thus, it produces more robust supervision than previous experts. It remains computationally manageable while preserving expressive power, which is why they are widely adopted in stochastic programming.
@@ 4.2 Expert Optimality
Even when employing the expert types defined above, the underlying optimisation problem may remain difficult to solve. Many deterministic optimisation problems are $\mathcal{NP}$-hard, making exact solution approaches computationally impractical in many real-world settings. As a result, practitioners often rely on heuristics or impose stopping criteria when using optimisation solvers. We therefore classify experts according to the optimality guarantees of the models they solve.
  • Task-Optimal Expert – A task-optimal expert solves its underlying optimisation model to optimality, assuming the problem is tractable and can be solved exactly within reasonable computational limits. Such experts are most applicable when efficient solution methods are available. For example, a strong branching expert evaluates linear programming relaxations for all fractional variables and selects the branching variable that yields the greatest objective improvement [Nair2020SolvingMI, alvarez2016online]. Optimal solutions may be obtained using specialised solvers, including mixed-integer programming solvers [sun2022learning], SAT solvers [sun2024enhancing], and dedicated solvers such as Concorde for the TSP [sun2021generalization]. We denote the policy of a task-optimal expert by $\pi^{*\cdot}$ and the corresponding optimal action by $a^{*\cdot}$, where $\cdot$ refers to the expert type in Table˜2.
  • Approximate Expert – An approximate expert produces actions that trade off solution quality for computational efficiency, aiming to generate near-optimal decisions. Common approximation strategies include:
    • Time limits: terminating the solver after a fixed time budget and returning the best solution found [li2025learning, baty2024combinatorial, pham2023prediction, sun2021generalization].
    • Optimality gaps: stopping once a predefined optimality gap is achieved [larsen2018predicting].
    • Objective-based refinement: applying alternative formulations or local improvements to progressively improve solution quality [ding2020accelerating].
    • Heuristics: directly using high-quality heuristic methods, such as hybrid genetic search for the prize-collecting TSP [baty2024combinatorial], or evaluating a finite set of candidate actions and selecting the best-performing one [song2020general].
@@ 4.4 Labelling strategy
      Single call - A single interaction with the expert yields multiple candidate actions for a given state. This occurs, for example, when solving an integer program using a mathematical programming solver that returns multiple feasible solutions. In [khalil2022mip], CPLEX is allowed to run for 60 minutes to generate a solution pool for each instance, terminating after identifying 1,000 solutions with objective values within 10% of the best solution found.
@@ 4.5 Number of Experts Used
  • Single Expert- The conventional setting in IL, where a single, typically optimal expert provides guidance. The learner assumes the expert is both reliable and inexpensive to query, consistently supplying accurate actions for all states. This simplifies learning but may limit adaptability in complex or diverse environments. Most literature uses a single expert to build the dataset.
  • Multi-Expert- In some cases, multiple experts with complementary specialisations are available. Their actions can be aggregated into a composite "meta-expert", combining diverse expertise to enhance robustness and flexibility across different aspects of the task. [song2020general] uses two different experts to get demonstrations. First, they use an expensive expert to gather initial expert demonstrations, which are used to train initial IL policies. Next, a simpler retrospective expert is used to augment the dataset. [rautenstrauss2025optimization] uses a set of non-optimal policies to label states for which the true hindsight expert is expensive. These states generally represent initial decision epochs in the horizon. [sun2023mega] introduces MEGA-DAgger, a variant of DAgger designed for interactive learning with multiple imperfect experts.
@@ 4.6 Trajectory Control
Moreover, to reduce reliance on costly expert calls, the expert is queried only for critical cases, such as novel or poorly handled states. This selective querying improves sample efficiency and focuses learning on high-impact scenarios. [menda2019ensembledagger] and [hoque2021lazydagger] describe the Ensemble DAgger and Lazy DAgger algorithms, where the expert is selectively queried based on the discrepancy measure between the expert and learned policy actions. However, such variants of the DAgger algorithms have not been experimented with for SDP in operations research.
@@
Using the proposed expert taxonomy, we introduce and evaluate an Adaptive DAgger-based algorithm that maps each dimension of the taxonomy to an explicit design choice in the learning procedure. This framework enables a structured comparison of expert configurations within an action imitation (AIT) learning setting.
@@ 5 Algorithm
In this section, we present a general-purpose adaptive algorithm for learning from experts in sequential decision problems. The proposed algorithm, shown in Algorithm˜3, explicitly incorporates the expert taxonomy introduced in Section˜4. A summary of the algorithm parameters is provided in Table˜3.
Table 3: List of parameters for Algorithm˜3
@@
The algorithm is based on the DAgger framework, which follows an iterative policy learning paradigm inspired by online learning and assumes access to one or more experts. Let $\mathcal{E}$ denote the set of available experts. At each learning iteration, indexed by $i$, the learner retrains its policy using all states encountered so far, together with the corresponding expert-labelled actions. This iterative process allows the learner to correct earlier mistakes and progressively improve their performance.
The procedure starts from an initial policy $\hat{\pi}_{0}$, which may be obtained from expert demonstrations or derived from a heuristic. This policy is then executed for $H$ simulated episodes using the decision rule, each consisting of $K$ decision epochs. At each epoch, the learner observes the current state and may query an expert up to $J$ times, as determined by the function $F_{j}$. Each query returns an expert action, potentially from different experts in $\mathcal{E}$, decided by the expert-selection function $F_{e}$.
We denote an expert action by $a^{\cdot}_{ijk}$, where $i$ indexes the learning iteration, $j$ the expert call within the epoch, and $k$ the decision epoch. For clarity, we omit the episode index in this notation. This may be task-optimal or approximate depending on the complexity of the underlying optimisation problem. The parameter $\kappa$ specifies the list of expert stopping criteria, thereby controlling the quality of the returned action. The resulting state–action pairs are stored in the dataset before the transition to the next state occurs.
After completing $H$ episodes, the dataset is augmented with the newly collected expert-labelled samples, and a new policy $\hat{\pi}_{i+1}$ is trained. This process is repeated until a stopping condition is satisfied, yielding progressively refined policies. The following paragraphs relate the components of the expert taxonomy to Algorithm˜3.
@@
4   Initialize dataset $\mathcal{D}\leftarrow\emptyset$; initial policy $\hat{\pi}_{0}$; iteration counter $i\leftarrow 0$;
5   learning routine $\textsc{Train}(\cdot)$;